Grade 12Differential Calculus

How function f(x) = (tan x) – x is always increasing function?

Siddharth verma
12 Years agoGrade 12

1 Answer

Jitender Singh
12 Years ago
Ans:
f(x) = tanx - x
f^{'}(x) = sec^{2}x - 1 = tan^{2}x \geq 0
So thats why f(x) is always increasing.
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty